# EdgeProfiler: A Fast Profiling Framework for Lightweight LLMs on Edge Using Analytical Model

### v2

@@ Abstract
This paper introduces EdgeProfiler, a fast profiling framework designed for evaluating lightweight Large Language Models (LLMs) on edge systems. While LLMs offer remarkable capabilities in natural language understanding and generation, their high computational, memory, and power requirements often confine them to cloud environments. EdgeProfiler addresses these challenges by providing a systematic methodology for assessing LLM performance in resource-constrained edge settings. The framework profiles compact LLMs, including TinyLLaMA, Gemma3.1B, Llama3.2-1B, and DeepSeek-r1-1.5B, using aggressive quantization techniques and strict memory constraints. Analytical modeling is used to estimate latency, FLOPs, and energy consumption. The profiling reveals that 4-bit quantization reduces model memory usage by approximately 60–70%, while maintaining accuracy within 2–5% of full-precision baselines. Inference speeds are observed to improve by 2–3$\times$ compared to FP16 baselines across various edge devices. Power modeling estimates a 35–50% reduction in energy consumption for INT4 configurations, enabling practical deployment on hardware such as Raspberry Pi 4/5 and Jetson Orin Nano Super. Our findings emphasize the importance of efficient profiling tailored to lightweight LLMs in edge environments, balancing accuracy, energy efficiency, and computational feasibility.
@@ I Introduction
Reducing dependency on cloud connectivity is also a significant advantage. Moving LLM inference to the edge allows applications to function independently of network conditions, ensuring availability even in areas with limited or no internet access. For example, a local LLM could manage user notes or provide contextual assistance directly on a smartphone, regardless of network strength. Moreover, local inference significantly reduces latency, which is crucial for real-time applications. Unlike cloud-based models whose response times are affected by network stability and speed, on-device inference ensures prompt interaction, enhancing the user experience. Fig. 1 illustrates the process where LLM training is performed in a cloud GPU cluster within a data center environment, and pre-trained LLM models are then deployed to edge devices, which independently handle local inference computation.
However, deploying LLMs on edge devices remains challenging due to their complex architectures and high memory and power requirements [2]. Models such as GPT-3 and LLaMA, with billions of parameters, are not feasible for resource-constrained environments [3, 4]. This has led to an active area of research focused on quantization techniques, model pruning, and efficient inference strategies that adapt LLMs to the limitations of edge hardware [5]. These efforts aim to strike a balance between computational feasibility, energy efficiency, and model accuracy.
Figure 1: LLM inference deployment on diverse edge devices, highlighting the critical need for accurate profiling to ensure efficient computation, reduced latency, and optimized energy consumption in resource-constrained setups.
In this paper, we introduce EdgeProfiler, a fast profiling framework designed to systematically evaluate the performance of lightweight LLMs in resource-constrained environments. The findings indicate that aggressive quantization significantly reduces the memory and computational requirements of lightweight LLMs, enabling practical deployment on resource-constrained edge devices. The slight impact on model accuracy is outweighed by improvements in memory efficiency and faster inference speeds. Reduced energy consumption from low-bit precision enhances feasibility for power-limited environments, such as mobile devices and microcontrollers. Careful quantization strategies and efficient hardware utilization bridge the gap between LLM capabilities and the limitations of edge systems, paving the way for broader and more practical applications. The main contributions of this paper are as follows:
@@
The structure of the paper is as follows: Section II presents a technical overview that highlights essential techniques applied to AI models for improving application-specific adaptations. Section III provides a detailed functional overview of the EdgeProfiler framework and the specific techniques employed in this study. Section IV discusses the evaluation results, while Section V offers a review of related work. Finally, Section VI summarizes the key insights and conclusions from the study.
@@ II Background
Efficient deployment of LLMs on edge devices requires a deep understanding of techniques that balance computational cost, memory efficiency, and model accuracy. This section provides the foundational concepts and methodologies essential for such deployments. Key strategies include quantization approaches that reduce model size and computation while maintaining acceptable accuracy levels, and advanced techniques like Quantization-Aware Training (QAT) that ensure models remain robust under aggressive compression. By understanding the trade-offs between different quantization schemes and how they interact with model structure and hardware constraints, this background forms the basis for evaluating and optimizing LLM performance in resource-constrained environments.
@@ II-A Quantization techniques for efficient LLM inference
II-A1 Symmetric Quantization
In symmetric quantization, zero in floating-point space is exactly representable by zero in integer space. The mapping function can be defined as:
@@ II-A1 Symmetric Quantization
where $x_{\text{float}}$ is the original floating-point value, $x_{\text{int}}$ is the quantized integer value, and $s$ is a positive scaling factor (common across the tensor). The dequantization process simply scales back:
@@
II-A2 Asymmetric Quantization
In asymmetric quantization, a nonzero offset is introduced to better handle data distributions that are not centered at zero. The quantization function becomes:
@@ II-A2 Asymmetric Quantization
Asymmetric quantization is generally used for activations, where the dynamic range shifts during inference. Symmetric quantization is highly hardware-efficient because it eliminates the need for offset additions during inference. It is particularly well-suited for weight matrices, where the value distributions are typically centered around zero. An advantage of symmetric quantization is that it requires less memory and simpler hardware logic[7, 8]. However, a key disadvantage is that it can result in a higher mean squared error (MSE) compared to asymmetric quantization, especially when the data distribution is symmetric. Asymmetric quantization provides better representational fidelity for non-centered data but introduces extra computational overhead due to the zero-point handling.
@@
II-A3 Per-Tensor Quantization
In per-tensor quantization, a single scale factor and, optionally, a single zero-point are used to quantize the entire tensor. The same quantization parameters are applied uniformly across all elements. similar to the equation 3. Per-tensor quantization offers the advantages of simple implementation and high efficiency, particularly for hardware accelerators that prefer uniform scaling across all elements. However, it also has notable disadvantages: it fails to capture the variance across different channels or feature maps, and it can lead to larger quantization errors if some channels have significantly wider or narrower value ranges compared to others[9].
II-A4 Per-Channel Quantization
In per-channel quantization, each channel of each output feature in a convolutional layer or each row in a matrix has its own scale and zero-point. This allows finer control over the quantization process.
@@ II-A4 Per-Channel Quantization
Quantization-Aware Training (QAT). Quantization-Aware Training is a technique where quantization effects are simulated during the training phase itself, allowing the model to adapt to low-precision representations. During QAT, both the forward and backward passes emulate quantized operations, typically using fake quantization functions that maintain gradient flow[7]. This enables the model to learn parameters that are more robust to quantization noise, resulting in significantly better accuracy compared to naive post-training quantization[11]. Formally, during QAT, the model optimizes the following objective:
@@
QAT often uses lower bit-widths 8-bit or 4-bit during simulation while keeping high-precision master weights for gradient updates. By introducing quantization noise during training, QAT yields models that maintain higher accuracy after deployment, especially important for aggressive compression regimes in edge and embedded systems[3]. Despite promising benefits, LLM quantization presents unique challenges, such as attention layers and softmax output are particularly sensitive to quantization noise. Even if inputs are quantized to 4-bit/8-bit, intermediate results often require higher precision to avoid numerical instability. Some input sequences containing rare or domain-specific tokens can significantly stress low-precision models[10, 12]. Hardware supports efficient low-bit general matrix-matrix multiplication (GEMM) for arbitrary bitwidths, necessitating custom CUDA kernels. Fig. 2 shows the Symmetric quantization technique.
@@
The paper [13] provides more insight into the numerous uses of LLM implementation in edge devices, in addition to quantization and compression strategies that makes this integration possible. These applications all emphasize the importance of utilizing LLMs locally for resource-constrained devices.
@@
III Methodology
Fig. 3 presents a high-level overview of our lightweight profiling framework, which evaluates the end-to-end performance and energy cost of serving transformer-based language models on a range of edge platforms. The framework operates through a modular workflow: it takes as input the model configuration, hardware specifications, and numeric precision settings; internally, it performs analytical modeling of key performance metrics such as parameter count, FLOPs, memory footprint, latency, and energy consumption; and finally, it outputs detailed performance estimates, enabling trade-off exploration across different device-model-precision combinations. This setup provides a scalable method for understanding and comparing LLM behavior on resource-constrained systems without requiring full hardware deployment.
III-A Performance Modeling Framework
At the core of our framework is the EdgeProfile, which takes three inputs: model configuration, hardware configuration, and precision configuration.
@@ III-A Performance Modeling Framework
Equation 7 captures all of the model’s weight parameters. The four projection matrices per layer plus the two feed-forward matrices, and the input/output embeddings.
@@
Equation 8 accounts for both attention projections query, key, value, and output projection(Q,K,V,O) and the two matmuls in each feed‐forward block, as well as LayerNorm and Softmax overhead.
@@
Equation 9 includes storing all weights ($P\cdot B$), the hidden‐state activations for the full sequence ($S\!H\!B$), and the cached key/value pairs across layers ($2L\!S\!H\!B$).
III-B Latency Breakdown
Using the counts above, we estimate each component of on-chip inference time.
@@ III-B Latency Breakdown
Equation 10 performs all the arithmetic at the device’s effective compute throughput.
Memory‐bound latency
@@
Equation 11 captures the delay reading/writing weights, activations, and cache from on-chip memory.
I/O, Host‐to‐Device, and Network Overheads
Each is computed by dividing the total weight bytes ($P\cdot B$) or shard bytes by the respective bandwidth and utilization factor:
@@
Equation 12 is the latency to load model weights from disk or flash storage. Equation 13 is the latency to copy weights over PCIe or NVLink into GPU memory. Equation 14 is the latency to exchange a shard of key/value cache across nodes. We also provide a fine‐grained breakdown of attention‐projection, key/value matmuls, MLP, layernorm, and softmax by dividing each operator’s raw FLOP count by effective compute throughput.
III-C Energy Modeling
Energy per token is estimated as the sum of compute and data‐movement costs:
$E=\mathrm{FLOPs/token}\,\times\,e_{\mathrm{flop}}\;+\;M\,\times\,e_{\mathrm{byte}},$ (15)
where $e_{\mathrm{flop}}$ and $e_{\mathrm{byte}}$ are measured in joules per FLOP and per byte, respectively.
This comprehensive modeling approach allows for systematic exploration of design trade-offs, helping identify bottlenecks and optimize configurations for efficient LLM deployment on edge systems.
@@ III-C Energy Modeling
FP16 2.2GB $\sim$3.13GB 1$\times$ baseline
TinyLlama INT8 1.2GB $\sim$2.25GB 1.86$\times$ minor
INT4 644MB $\sim$1.78GB 2.45$\times$ moderate
FP16 2.0GB $\sim$2.44GB 1$\times$ baseline
Gemma3-1B INT8 1.1GB $\sim$1.60GB 1.26$\times$ minor
INT4 815MB $\sim$1.35GB 1.52$\times$ moderate
FP16 2.5GB $\sim$3.58GB 1$\times$ baseline
Llama3.2-1B INT8 1.3GB $\sim$2.53GB 2.7$\times$ minor
INT4 776MB $\sim$2.01GB 3.33$\times$ moderate
FP16 3.6GB $\sim$3.91GB 1$\times$ baseline
DeepSeek-r1-1.5B INT8 1.9GB $\sim$2.55GB 2.19$\times$ minor
INT4 1.1GB $\sim$1.84GB 2.97$\times$ moderate
IV Evaluation
This section evaluates the performance of lightweight LLMs on representative edge devices. The profiling framework is used to analyze trade-offs between model size, precision levels, and hardware characteristics. Key metrics such as latency, memory footprint, arithmetic intensity, and energy consumption are examined, offering insights into the feasibility and efficiency of deploying LLMs on resource-constrained platforms. The framework was executed on a standard workstation equipped with a 10th Gen Intel(R) Core(TM) i7-10700F CPU, 32 GB DDR4 memory, and without using a dedicated GPU, running Ubuntu 22.04 LTS.EdgeProfiler is available on GitHub: EdgeProfiler
@@ IV Evaluation
Precision reduction from FP32 to FP16 halves each component’s latency, and INT8 cuts it roughly by four. On Raspberry Pi 4, end-to-end latency drops from $\sim$15.4 s (FP32) to $\sim$3.9s (INT8), driven almost entirely by shorter I/O and transfer times of smaller weight footprints. However, even at INT8, I/O remains the bottleneck (3.5s vs. compute  0.13s), underscoring that the network and storage subsystems must be improved in tandem with quantization to realize truly low-latency inference shown in all the Fig. 4.
@@
The Jetson Nano Super’s higher storage bandwidth and PCIe host-to-device throughput, shown in Fig. 4(c), deliver a dramatic reduction in I/O cost: INT8 inference completes in $\sim$1.05s end-to-end, nearly four times faster than on the Raspberry Pi 5. Compute and memory latencies on the Jetson ($\sim$0.07s and $\sim$0.88s at FP32) are comparable to—though still lower than those on the Pis, but the real advantage comes from overlapping and hiding I/O behind faster transfers. This highlights that mid-range AI accelerators can shift the bottleneck away from storage if paired with efficient weight-delivery mechanisms.
Across all models and platforms, arithmetic intensity remains low (well under 1 FLOP/byte), confirming a data-movement–bound regime. As models grow, compute time will rise, but so too will weight volume, so similar bottlenecks are likely unless novel compression or on-device caching techniques reduce the need for repeated weight fetches. In practice, combining aggressive quantization with persistent in-RAM weight formats, alongside higher-bandwidth storage paths (e.g., NVMe, LPDDR5X), appears essential for sub-second LLM inference at the billion-parameter scale.
Ablation Studies on Quantization. Table II summarizes how reducing precision from FP16 to INT8 and INT4 impacts model size, peak memory usage, inference throughput, and observed accuracy degradation across four 1-1.5B (TinyLlama‐1B, Gemma3‐1B, Llama3.2‐1B, DeepSeek‐r1‐1.5B) and peak runtime memory (3.13GB to 2.25GB), while boosting inference speed by roughly 1.8×1.8$\times$1.8 × with only a minor accuracy loss. Further quantization to INT4 yields even more dramatic storage saving (down to 644MB for TinyLlama) and a 2.45×2.45$\times$2.45 × speedup, at thecost of a moderate drop in task performance,
@@
Across all architectures, INT8 delivers near-2×2$\times$2 × speed gain and $\sim$50reduction in memory footprint with negligible impact on accuracy. The large models Llama3.2-1B and DeepSeek-r1-1.5B benefit even more from low-precision storage, up to $\times3$ speedup, incur more noticeable accuracy degradation, suggesting their suitability only when memory and latency constraints are extreme and a moderate loss in quality is acceptable.
V Related Works
Recent research has focused on adapting LLMs for edge devices, addressing challenges like limited memory, compute capacity, and energy constraints. Work [14] empirically studies how resource-constrained computing environments affect the design choices for personalized LLMs, uncovering key trade-offs and offering practical guidelines for customization and deployment on edge devices. However, the empirical study might be limited to specific models and hardware, affecting generalizability. Deeploy [15] introduces a compiler framework that translates pre-trained Foundation Models into compact Small Language Models (SLMs) for microcontrollers, showcasing a practical approach for edge NLP tasks. Post-training quantization techniques, such as GPTQ [16], compress model weights to as low as 3–4 bits, balancing memory efficiency and accuracy. Complementary methods like AWQ [17] and SmoothQuant [18] further enhance memory and performance trade-offs.
MobileLLM [19] demonstrates that sub-billion parameter models can provide effective NLP functionality on mobile and edge platforms, achieving a balance between resource constraints and utility. BioMistral-7B [20] applies quantization and model merging to tailor LLMs for biomedical tasks, combining compression with domain-specific optimization for constrained environments. Memory innovations, exemplified by “LLM in a Flash” [21], reduce DRAM usage by dynamically loading only essential model weights from flash storage during inference. This approach addresses limitations of static weight loading, improving LLM feasibility for edge deployment. Another recent advancement in efficient LLM deployment is the work on automatic INT4 weight-only quantization and optimized CPU inference, as described by the authors in [22]. This approach introduces an automatic quantization flow and a specialized LLM runtime with highly optimized kernels, enabling extreme inference efficiency on CPUs. Their method is demonstrated on popular models including LLaMA2, LLaMA, and GPT-NeoX, highlighting the potential for practical, low-precision LLM deployment on standard hardware. This complements efforts like GPTQ and “LLM in a Flash” by extending efficient deployment strategies to CPUs, further broadening the applicability of LLMs in resource-constrained environments.
The works discussed collectively highlight various strategies such as compact model generation, quantization, and innovative memory management are necessary to deploy LLMs on resource-constrained edge devices. This underscores the pressing need for a comprehensive profiling framework specifically designed for edge environments. To the best of our knowledge, no existing approach offers an analytical modeling solution for this challenge. EdgeProfiler addresses this gap by taking both the model and hardware configuration as inputs, while targeting a desired level of precision in the performance evaluation. It does not involve architecture level simulators such as gem5 [23] rather a set of analytical models which makes it faster profiling and is easier to implement.

### v3

@@ Abstract
[…]
@@ I Introduction
Reducing dependency on cloud connectivity is also a significant advantage. Moving LLM inference to the edge allows applications to function independently of network conditions, ensuring availability even in areas with limited or no internet access. For example, a local LLM could manage user notes or provide contextual assistance directly on a smartphone, regardless of network strength. Moreover, local inference significantly reduces latency, which is crucial for real-time applications. Unlike cloud-based models whose response times are affected by network stability and speed, on-device inference ensures prompt interaction, enhancing the user experience. However, deploying LLMs on edge devices remains challenging due to their complex architectures and high memory and power requirements [2]. Models such as GPT-3 and LLaMA, with billions of parameters, are not feasible for resource-constrained environments [3, 4]. This has led to an active area of research focused on quantization techniques, model pruning, and efficient inference strategies that adapt LLMs to the limitations of edge hardware [5]. These efforts aim to strike a balance between computational feasibility, energy efficiency, and model accuracy.
In this paper, we introduce EdgeProfiler, a fast profiling framework designed to systematically evaluate the performance of lightweight LLMs in resource-constrained environments. It shows that aggressive quantization makes lightweight LLMs practical for edge deployment by reducing memory, computation, and energy costs with only minor accuracy trade-offs. The main contributions of this paper are as follows:
@@ II-A Quantization techniques for efficient LLM inference
Symmetric Quantization: In symmetric quantization, zero in floating-point space is exactly representable by zero in integer space. The mapping function can be defined as:
@@
where $x_{\text{float}}$ is the original floating-point value, $x_{\text{int}}$ is the quantized integer value, and $s$ is a common positive scaling factor. Dequantization process scales back as:
@@
Asymmetric Quantization: Asymmetric quantization adds a nonzero offset to handle data distributions not centered at zero, making the quantization function:
@@
Asymmetric quantization is generally used for activations, where the dynamic range shifts during inference. Symmetric quantization is highly hardware-efficient because it eliminates the need for offset additions during inference. It is particularly well-suited for weight matrices, where the value distributions are typically centered around zero. An advantage of symmetric quantization is that it requires less memory and simpler hardware logic[7, 8]. However, a key disadvantage is that it can result in a higher mean squared error (MSE) compared to asymmetric quantization.
@@
Per-Tensor Quantization: In per-tensor quantization, a single scale factor and, optionally, a single zero-point are used to quantize the entire tensor. The same quantization parameters are applied uniformly across all elements. similar to the equation 3. Per-tensor quantization offers the advantages of simple implementation and high efficiency, particularly for hardware accelerators that prefer uniform scaling across all elements. However, it also has notable disadvantages: it fails to capture the variance across different channels or feature maps, and it can lead to larger quantization errors if some channels have significantly wider or narrower value ranges compared to others[9].
Per-Channel Quantization: In per-channel quantization, each channel of each output feature in a convolutional layer or each row in a matrix has its own scale and zero-point. This allows finer control over the quantization process.
@@
Quantization-Aware Training (QAT). Quantization-Aware Training is a technique where quantization effects are simulated during the training phase itself, allowing the model to adapt to low-precision representations. During QAT, both the forward and backward passes emulate quantized operations, typically using fake quantization functions that maintain gradient flow[7]. This enables the model to learn parameters that are more robust to quantization noise, resulting in significantly better accuracy compared to naive post-training quantization[11]. During QAT, the model optimizes following objective:
@@
QAT often uses lower bit-widths 8-bit or 4-bit during simulation while keeping high-precision master weights for gradient updates. By introducing quantization noise during training, QAT yields models that maintain higher accuracy after deployment, especially important for aggressive compression regimes in edge and embedded systems[3]. Despite promising benefits, LLM quantization presents unique challenges, such as attention layers and softmax output are particularly sensitive to quantization noise. Even if inputs are quantized to 4-bit/8-bit, intermediate results often require higher precision to avoid numerical instability. Some input sequences containing rare or domain-specific tokens can significantly stress low-precision models[10, 12, 13]. Fig. 1 shows the Symmetric quantization technique.
@@
III Related Works
Lightweight LLMs. MobileLLM [19] demonstrates that sub-billion parameter models can provide effective NLP functionality on mobile and edge platforms, achieving a balance between resource constraints and utility. BioMistral-7B [20] applies quantization and model merging to tailor LLMs for biomedical tasks, combining compression with domain-specific optimization for constrained environments. Memory innovations, exemplified by “LLM in a Flash” [21], reduce DRAM usage by dynamically loading only essential model weights from flash storage during inference. This approach addresses limitations of static weight loading, improving LLM feasibility for edge deployment. Another recent advancement in efficient LLM deployment is the work on automatic INT4 weight-only quantization and optimized CPU inference, as described by the authors in [22].
LLM Inference on Edge Devices. Recent research has focused on adapting LLMs for edge devices, addressing challenges like limited memory, compute capacity, and energy constraints. Work [14] empirically studies how resource-constrained computing environments affect the design choices for personalized LLMs, uncovering key trade-offs and offering practical guidelines for customization and deployment on edge devices. However, the empirical study might be limited to specific models and hardware, affecting generalizability. Deeploy [15] introduces a compiler framework that translates pre-trained Foundation Models into compact Small Language Models (SLMs) for microcontrollers, showcasing a practical approach for edge NLP tasks. Post-training quantization techniques, such as GPTQ  [12], compress model weights to as low as 2–4 bits, balancing memory efficiency and accuracy. Complementary methods like AWQ [17] and SmoothQuant [18] further enhance memory and performance trade-offs.
IV Methodology
IV-A Performance Modeling Framework
At the core of our framework is the EdgeProfile, which takes three inputs: model configuration, hardware configuration, and precision configuration, as shown in the figure. 2.
@@ IV-A Performance Modeling Framework
The total number of model parameters is fundamental for determining storage, initialization time, and memory footprint. EdgeProfiler calculates it by summing contributions from all projection and embedding matrices across the transformer architecture:
@@
Equation 7 computes the total number of model parameters by summing contributions from self-attention, feed-forward, and embedding layers. The first term, $L(4H^{2})$, counts four linear projection matrices (query, key, value, output) per layer, each of size $H\times H$. The second term, $L(2HI)$, counts the two linear layers in each feed-forward network with input dimension $H$ and expansion dimension $I$. The final term, $2VH$, accounts for input and output embedding matrices that project vocabulary tokens to hidden representations and back. This formulation allows EdgeProfiler to accurately capture total model size, critical for estimating storage requirements, weight loading times, and parameter transfer bandwidth needs during initialization and deployment.
@@
To estimate the compute workload per token, EdgeProfiler calculates the total floating-point operations required for processing:
@@
Equation 8 calculates the total floating-point operations required to process a single token. The $6H^{2}$ term aggregates four projections for attention and two residual projections, while $4HS$ models token-wise dot-product attention computation over sequence length $S$. The terms $4HI+4IH$ capture feed-forward block matmuls, and $9H$ accounts for LayerNorm, bias additions, and other elementwise operations. This equation enables precise estimation of compute workload per token, which is essential for understanding whether inference is compute-bound or memory-bound on a target edge device.
@@
Estimating total memory usage is essential for ensuring inference fits within device capacity:
@@
Equation 9 estimates total memory usage, combining model weights ($P\cdot B$), hidden-state activations ($S\cdot H\cdot B$), and cached key/value tensors ($2L\cdot S\cdot H\cdot B$). Here, $B$ is bytes per parameter. The key/value cache term reflects storage of projected keys and values for each token in each layer, enabling efficient autoregressive decoding. This memory estimation is vital for determining whether an edge GPU or accelerator has sufficient on-chip SRAM or VRAM to store the model without offloading to external DRAM, which incurs additional latency and energy. Accurate memory modeling guides system designers in choosing appropriate batch sizes, quantization schemes, and layer fusion strategies to fit within device constraints.
IV-B Latency Breakdown
@@
EdgeProfiler estimates compute-bound inference time as:
@@
Equation 10 estimates inference time under compute-bound conditions by dividing total FLOPs per token by the effective device compute throughput. $\mathrm{peak}_{\mathrm{flops}}$ is the advertised peak performance of the accelerator, and $U_{\mathrm{compute}}$ is a utilization factor reflecting kernel launch overheads, pipeline stalls, and non-ideal compute scheduling. This latency model helps determine if performance is limited by computational capacity or by data transfer bottlenecks. By adjusting $U_{\mathrm{compute}}$, users can model performance improvements from kernel fusion, operator scheduling, or optimized hardware utilization techniques such as Tensor Cores or systolic arrays in edge accelerators.
Memory latency
Memory latency computed as:
@@
Equation 11 models the time to read or write total data volume $M$ across the device’s memory subsystem. $\mathrm{memory}_{\mathrm{BW}}$ is the theoretical bandwidth, and $U_{\mathrm{memory}}$ is utilization efficiency capturing factors such as bank conflicts, non-coalesced accesses, and DRAM refresh penalties. This estimation is essential for memory-bound workloads, where data movement rather than arithmetic dominates runtime. It guides decisions on operator reordering, memory tiling, quantization to reduce data size, and architectural choices such as increasing on-chip SRAM to reduce external DRAM usage.
I/O, Host-to-Device, and Network Overheads
Estimates of data transfer overheads in critical edge deployments:
@@
Equations 12-14 estimate data transfer overheads. Equation 12 models loading weights from storage (e.g. SSD or flash), 13 models copying weights to device memory over PCIe or NVLink, and 14 models exchanging key/value cache shards over a network in distributed settings. Each uses bandwidth and utilization terms to convert bytes to latency in milliseconds. These overheads are critical for edge systems where PCIe Gen3/Gen4 or embedded interconnect bandwidths can become bottlenecks.
IV-C Energy Modeling
EdgeProfiler models energy consumption by Total energy per token as:
$E=\mathrm{FLOPs/token}\,\times\,e_{\mathrm{flop}}\;+\;M\,\times\,e_{\mathrm{byte}}$ (15)
Equation 15 computes token-level energy as the sum of compute energy (FLOPs $\times$ energy per FLOP) and memory energy (total bytes moved $\times$ energy per byte). $e_{\mathrm{flop}}$ and $e_{\mathrm{byte}}$ are hardware-specific constants obtained from power modeling or direct measurement. This metric is critical for battery-constrained edge deployment, enabling rapid estimation of average and peak energy consumption per inference.
@@ IV-C Energy Modeling
[…]
V Evaluation
This section evaluates the performance of lightweight LLMs on representative edge devices. The profiling framework is used to analyze trade-offs between model size, precision levels, and hardware characteristics. Key metrics such as latency, memory footprint, arithmetic intensity, and energy consumption are examined. The framework was executed on a standard workstation equipped with a 10th Gen Intel(R) Core(TM) i7-10700F CPU, 32 GB DDR4 memory, and without using a dedicated GPU, running Ubuntu 22.04 LTS. EdgeProfiler is available on GitHub:EdgeProfiler
@@ V Evaluation
Precision reduction from FP32 to FP16 halves each component’s latency, and INT8 cuts it roughly by four. On Raspberry Pi 4, end-to-end latency drops from $\sim$15.4 s (FP32) to $\sim$3.9s (INT8), driven almost entirely by shorter I/O and transfer times of smaller weight footprints. However, even at INT8, I/O remains the bottleneck (3.5s vs. compute  0.13s), underscoring that the network and storage subsystems must be improved in tandem with quantization to realize truly low-latency inference shown in all the Fig. 3. The Jetson Nano Super’s higher storage bandwidth and PCIe host-to-device throughput, shown in Fig. 3(c), deliver a dramatic reduction in I/O cost: INT8 inference completes in $\sim$1.05s end-to-end, nearly four times faster than on the Raspberry Pi 5. Compute and memory latencies on the Jetson ($\sim$0.07s and $\sim$0.88s at FP32) are comparable to—though still lower than those on the Pis, but the real advantage comes from overlapping and hiding I/O behind faster transfers. This highlights that mid-range AI accelerators can shift the bottleneck away from storage if paired with efficient weight-delivery mechanisms.
@@
Ablation Studies on Quantization. Table II summarizes that reducing precision from FP16 to INT8 and INT4 impacts model size, peak memory usage, inference throughput. We observe accuracy degradation across four 1-1.5B (TinyLlama‐1B, Gemma3‐1B, Llama3.2‐1B, DeepSeek‐r1‐1.5B) and peak runtime memory (3.13GB to 2.25GB) but a boost in inference speed by roughly $1.8\times$ with only a minor accuracy loss. Further quantization to INT4 yields even more storage saving (644MB for TinyLlama) and a $2.45\times$ speedup, at the cost of a moderate drop in task performance. Across all architectures, INT8 delivers near-$2\times$ speed gain and $\sim$50% reduction in memory footprint with negligible impact on accuracy. The large models Llama3.2-1B and DeepSeek-r1-1.5B benefit even more from low-precision storage, up to $3.3\times$ speedup, incur more noticeable accuracy degradation, suggesting their suitability only when memory and latency constraints are extreme and a moderate loss in quality is acceptable.